# How Breakable Is Privacy: Probing and Resisting Model Inversion Attacks in Collaborative Inference

### v4

@@ 3.1. Knowledge background of adversary
The gray-box scenario enables the adversary to leverage their partial understanding of the edge model’s architecture to enhance the capability of their attack model. For instance, if the defender partitions the first three convolutional layers of a ResNet-18 model (He et al., 2016) and further deepens the network, while the adversary is only aware of the original ResNet-18 architecture, they can only construct an approximate attack model based on partial knowledge of the initial three layers. However, adversaries can address this limitation in this scenario by utilizing their understanding of the edge model’s architecture and reasoning about its mapping capabilities to refine their attack model.
@@
This study focuses on image classification but emphasizes that the defense methods apply to other tasks where the edge model primarily handles feature extraction. Since segmentation, detection, and classification also rely on edge-processed features, MIA and defenses are task-agnostic. Prior work (Azizian and Bajić, 2024) has demonstrated input reconstruction and defense effectiveness in object detection using Gen-based MIA.
Moreover, image data is particularly vulnerable to MIAs due to its high dimensionality, as shallow convolutional layers retain more redundant information (Zhang et al., 2020; He et al., 2016). This increases the challenge of protecting image data, making it an ideal test case for privacy defenses. In contrast, textual data, with its strong contextual dependencies and syntactic complexity, has shown limited success in achieving high-quality MIA reconstructions (Dibbo, 2023; Zhang et al., 2022). This study focuses on image data due to its interpretability, making it suitable for evaluating privacy defenses.
3.2. Attack strategy of the adversary
In the white-box attack scenario, the adversary can perform attacks using the MLE-based MIA method proposed in (He et al., 2019, 2020). To ensure the experiments are both comparative and representative, hyper-parameter adjustments were made to achieve performance closer to optimal.
In the black-box attack scenario, the adversary can employ methods from (Yang et al., 2019; He et al., 2019; Yin et al., 2023), relying solely on output features for reverse mapping without access to the edge model’s internal details. When protected by NND & IB-based defenses, the adversary lacks knowledge of architectural changes and must design the attack model based on prior assumptions.
@@ 3.2. Attack strategy of the adversary
4. Theoretical Analysis of MIA in CI
4.1. Principles analysis of MIA
@@
(3) $\begin{gathered}x^{(k+1)}=x^{(k)}-\eta\nabla_{x}\mathcal{L}\left(x,f_{\text{edge }}\left(x^{(k)}\right),z\right),\\ x^{*}=\arg\min_{x}\mathcal{L}\left(x,f_{\text{edge }}\left(x^{(k)}\right),z\right)\end{gathered}$
@@
According to (He et al., 2019), the loss function combines the $l_{2}$ distance between $f_{\text{edge }}\left(x^{(k)}\right)$ and $z$ with a TV regularization term for $x$, forming the basis of rMLE-based MIA. The method relies on gradient descent, where the loss values depend on output features, and gradients are computed using the target model’s parameters.
To achieve effective attacks, sufficient output feature information $\delta(z)$ and a well-chosen loss function are essential to ensure optimization space. Additionally, white-box knowledge of the edge model is required, as complex model mappings and limited feature information increase the risk of gradient descent converging to local minima (Struppek et al., 2022; Zhang et al., 2020). For instance, if two similar images produce nearly indistinguishable output features, reconstruction becomes challenging.
When dealing with complex edge model mappings or models protected by defense techniques, directly computing gradients in MLE-based MIA can significantly hinder optimization performance. To address this, He et al. (He et al., 2019, 2020), building on the method from (Yang et al., 2019), developed a reverse mapping network for the edge model, which is categorized as Gen-based MIA. The key idea is to use a generator to approximate the reverse mapping of the edge model. Different methods adopt various strategies to train the generator. For instance, Yang et al. (Yang et al., 2019) trained a decoder using auxiliary task-related datasets, while Yin et al. (Yin et al., 2023) utilized intercepted feature data for training. The principle of this attack is summarized as follows.
(4) $\begin{gathered}G=\arg\min_{G}\mathbb{E}_{x^{\prime}\sim p\left(x^{\prime}\right)}\mathcal{L}\left(G\left(f_{\text{edge }}\left(x^{\prime}\right)\right),x^{\prime}\right),\\ x^{*}=G(z)\end{gathered}$
@@
  1. (1)
    A well-matched training dataset aligned with the target task, ideally with a distribution similar to the target dataset (Yang et al., 2019).
  2. (2)
@@
The rationale is that if the information in $f_{\text{edge }}\left(x^{\prime}\right)$ is insufficient or cannot independently represent $x^{\prime}$’s features, the loss in (4) will struggle to converge as $G\left(f_{\text{edge }}\left(x^{\prime}\right)\right)$ remains ambiguous, limiting its training. Among them, the maximum separability is more decisive. For instance, even with random feature drops, if the remaining features distinguish inputs well, the attack can still reconstruct effectively (He et al., 2020). Such frameworks resemble DAE (Gondara, 2016), where the decoder can adapt to noise or feature drops.
@@ 4.2. Difficulty criterion for MIA in CI
These methods primarily compare $z$ with generated features, but a poorly trained or protected model can produce $z$ with high $\delta(z)$ yet low $H(z)$, indicating suboptimal separability. This weakens attack effectiveness despite the apparent information sufficiency. Thus, MLE-based MIA difficulty is more influenced by $I(x;z)$, $H(x|z)$, and feature entropy $H(z)$.
We can demonstrate the influence of mutual information and entropy on MLE-based MIA using the following relationship for any $x\to z\to x^{*}$, derived from Fano’s inequality (Arrow, 1969; Xu et al., 2025):
(6) $H\left(P_{e}\right)+P_{e}\log|\mathbb{X}|\geq H(x\mid x^{*})\geq H(x\mid z)$
where $P_{e}=Prob\{x\neq x^{*}\}$, $x^{*}$ is the reconstruction by MIA, and $\mathbb{X}$ denotes the value space of $x$.
Since $\mathbb{X}\geq 2$ and $H\left(P_{e}\right)\leq 1$. Therefore, inequality (6) can be written as
(7) $P_{e}\geq\frac{H(x\mid z)-1}{\log|\mathbb{X}|}=\frac{H(x)-I(x,z)-1}{\log|\mathbb{X}|}=\frac{H(x)-H(z)-1}{\log|\mathbb{X}|}$
Since $H(x)$ and $|\mathbb{X}|$ are influenced by training data, they are typically assumed to be objective and invariant. Therefore, we observe that increasing the lower bound of $P_{e}$, and thus the difficulty of MIA, is influenced by $H(x\mid z)$, $H(z)$, and $I(x,z)$.
Factors influencing Gen-based MIA. From Equation (4), Gen-based MIA requires sufficient input information $\delta(z)$ and high $H(z)$ to ensure maximum separability and enable loss convergence, allowing the generator to simulate reverse mapping effectively. A larger $\delta(z)$ not only facilitates generator training but also directly impacts the reconstruction of $x^{*}$. At the same time, the relationship in Inequality (7) also holds here (Zhang et al., 2023). Therefore, in addition to the factors influencing MLE-based MIA, Gen-based MIA is further effectively affected by $\delta(z)$.
@@
(8) $D_{\text{mia}}\propto\left(\frac{k_{1}}{\{I(x;z);H(x\mid z);H(z)\}}+\frac{k_{2}}{\delta(z)}\right)$
Based on this criterion, existing defenses (Wang et al., 2021; Peng et al., 2022; Duan et al., 2023) using the information bottleneck theory can be understood as approximating a reduction in $I(x;z)$ through various algorithms to increase the difficulty of MIA. Perturbation-based defenses (He et al., 2020) aim to directly increase $H(x\mid z)$. The defensive effect of deeper neural networks (Ding et al., 2024; Azizian and Bajić, 2024) can be explained by their ability to reduce $I(x;z)$ and $\delta(z)$. This is mainly because, without skip connections (Hao et al., 2025), mutual information progressively decreases during forward propagation in neural networks. In contrast, residual networks with skip connections preserve or even enhance $\delta(z)$ by adding layer outputs to inputs, but the decline in mutual information still poses challenges for MIA (see section 4.3 for experimental cases).
4.3. Demonstration and Guidance for Defense
To validate $D_{mia}$ and the theoretical analysis, we designed experiments using a CNN that progressively reduces the spatial size while increasing the channel size, based on the framework in (Yang et al., 2019; He et al., 2016), and ResNet-18 with skip connections, quantifying each block. We trained MINE (Belghazi et al., 2018) on the same dataset used to train the full model $f$ to estimate the lower bound of $I(x,z)$. The core idea of MINE is to optimize a neural network estimator using the Donsker-Varadhan (DV) representation (Donsker and Varadhan, 1983) as the loss function, providing an approximate estimate of the mutual information between random variables.
@@ 4.3. Demonstration and Guidance for Defense
(9) $I(X;Z)=\mathbb{E}_{p(x,z)}\left[\log\frac{p(x,z)}{p(x)p(z)}\right]$
@@
(10) $I(X;Z)\geq\sup_{T}\mathbb{E}_{p(x,z)}[T(x,z)]-\log\mathbb{E}_{p(x)p(z)}\left[e^{T(x,z)}\right]$
@@
(11) $\hat{I}(X;Z)=\mathbb{E}_{p(x,z)}\left[T_{\theta}(x,z)\right]-\log\mathbb{E}_{p(x)p(z)}\left[e^{T_{\theta}(x,z)}\right]$
As shown in Figure 2, we trained the target models on the CIFAR-10 (Krizhevsky et al., 2009) and used Equation (11) to train MINE for the corresponding blocks. Figures 2(a, b) show that increasing network depth reduces both mutual information and effective information, leading to a decrease in attack performance. Figures 2(c, d) illustrate that while skip connections within two layers do not reduce $\delta(z)$, they slightly lower $I(x,z)$, moderately affecting attack performance. Previous research (Hao et al., 2025) identified vulnerabilities caused by skip connections in centralized MIA. Our study further highlights their impact on MIA effectiveness in CI.
Figure 2. The impact of neural network depth on $I(x,z)$, $\delta(z)$, as well as MLE & Gen-based MIA.
[…]
2
@@
Figure 3. Overview of SiftFunnel. Specifically, the design is divided into two aspects: the architecture and the loss function for the edge and cloud model.
[…]
3
Notably, while mutual information at ResNet’s fourth layer is higher than that at CNN’s second layer, MLE-based MIA does not perform better. This stems from the fact that these methods rely heavily on gradient descent, making them sensitive to suboptimal optimizers, hyper-parameter settings, and deeper network structures, which hinder gradient calculations for optimizing $x$. The lack of flexible and efficient methods for MLE-based MIA remains a gap in current research, often leading to their exclusion in defense evaluations—a topic warranting further exploration (Ding et al., 2024). In our experiment and Section 6, we still evaluate MLE-based MIA against various defenses and provide evidence demonstrating the effectiveness of our method in defending against such attacks.
@@
5. SiftFunnel: Information Sifting Funnel
5.1. Overview of SiftFunnel
@@
Specifically, in the edge model design, the final layer adopts a funnel-shaped structure, reducing feature channels while preserving spatial dimensions. This reduces computation, lightens the edge device load, and limits redundant information. Attention mechanisms and multi-receptive field convolutions enhance task-relevant information, while a cloud-based upsampling module preserves accuracy. To reduce redundancy, $I(x,z)$ and increase the uncertainty of $z$ to $x$, we use Distance Correlation (Li et al., 2012) for nonlinear correlations and the Pearson Coefficient (Cohen et al., 2009) with $l_{1}$-norm for linear correlations. Additionally, Label Smoothing (LS) and Kullback-Leibler (KL) divergence (Donsker and Varadhan, 1983) are used during training to improve model usability while reducing the separability of representations in $z$.
5.2. The Design of Shallow Edge Model
@@
Information Funnel. Deepening the neural network enhances feature extraction and reduces $I(x,z)$, as confirmed in Section 4. Previous work (Ho et al., 2024) shows through Fisher information (Rissanen, 1996) that deeper networks improve task relevance and resist MIA. Inspired by this, we redesign the final layer of the edge model with an Information Funnel module. This module reduces output channels to one-fourth in the first layer and further in the second layer, with channel selection based on the spatial size of feature maps. If the spatial size is reduced significantly, channels can exceed 10 to maintain accuracy. If unchanged, channels can be set to 2 to limit redundancy while preserving accuracy. Adjusting channels based on spatial size is necessary because spatial information loss is irreversible, whereas channel information loss can be compensated in the cloud model. This compensation increases $I(z,y)$ without reversing the reduction in $I(x,z)$, maintaining high attack difficulty as shown in Section 6.3. So, we add the channel compensation module in the cloud model to guarantee usability. Importantly, assumptions about extracting features from non-independent transmission modules are idealized, requiring adversaries to have full white-box knowledge of the cloud model as if they were the developer.
@@
The Information Funnel’s channel structure controls output but does not limit redundancy or enhance task-relevant information transmission, affecting usability. Inspired by Inception networks (Szegedy et al., 2017), we redesign it as a multi-receptive field convolutional framework with three low-channel layers. A $1*1$ kernel fuses channel information, followed by $3*3$ and $5*5$ kernels to expand the receptive field, extract richer task-relevant features, improve decision-making, and reduce mutual information with $x$, while lowering parameters of the edge model and controlling $\delta(z)$.
@@
Attention Block. To enhance the output of task-relevant information from the Information Funnel, attention modules are incorporated both before and after it. The first block combines an SE block (Hu et al., 2018) and Triplet Attention (Misra et al., 2021). The SE block evaluates the contribution of information within each channel, suppressing channels with low task relevance. Triplet Attention reduces spatial information while maintaining smooth, continuous outputs, preserving fine-grained features for further processing. The second block integrates SE and CBAM modules (Woo et al., 2018). Compared to Triplet Attention, CBAM is a lightweight module that sequentially calculates channel and spatial weights to enhance or suppress input features. Since CBAM’s channel and spatial attention are independently designed, certain weights are suppressed redundantly, resulting in sparser outputs. This sparsity is beneficial for output processing, effectively increasing $H(x\mid z)$ while reducing $H(z)$ and $\delta(z)$.
5.3. The Design of Loss Function

### v5

@@ 3.1 Knowledge background of adversary
The gray-box scenario enables the adversary to leverage their partial understanding of the edge model’s architecture to enhance the capability of their attack model. For instance, if the defender partitions the first three convolutional layers of a ResNet-18 model [38] and further deepens the network, while the adversary is only aware of the original ResNet-18 architecture, they can only construct an approximate attack model based on partial knowledge of the initial three layers. However, adversaries can address this limitation in this scenario by utilizing their understanding of the edge model’s architecture and reasoning about its mapping capabilities to refine their attack model.
@@
This study focuses on image classification but emphasizes that the defense methods apply to other tasks where the edge model primarily handles feature extraction. Since segmentation, detection, and classification also rely on edge-processed features, MIA and defenses are task-agnostic. Prior work [15] has demonstrated input reconstruction and defense effectiveness in object detection using Gen-based MIA.
Moreover, image data is particularly vulnerable to MIAs due to its high dimensionality, as shallow convolutional layers retain more redundant information [11, 38]. This increases the challenge of protecting image data, making it an ideal test case for privacy defenses. In contrast, textual data, with its strong contextual dependencies and syntactic complexity, has shown limited success in achieving high-quality MIA reconstructions [39, 40]. This study focuses on image data due to its interpretability, making it suitable for evaluating privacy defenses.
3.2 Attack strategy of the adversary
In the white-box attack scenario, the adversary can perform attacks using the MLE-based MIA method proposed in [6, 16]. To ensure the experiments are both comparative and representative, hyper-parameter adjustments were made to achieve performance closer to optimal.
In the black-box attack scenario, the adversary can employ methods from [10, 6, 8], relying solely on output features for reverse mapping without access to the edge model’s internal details. When protected by NND & IB-based defenses, the adversary lacks knowledge of architectural changes and must design the attack model based on prior assumptions.
@@ 3.2 Attack strategy of the adversary
4 Theoretical Analysis of MIA in CI
4.1 Principles analysis of MIA
@@
$\begin{gathered}x^{(k+1)}=x^{(k)}-\eta\nabla_{x}\mathcal{L}\left(x,f_{\text{edge }}\left(x^{(k)}\right),z\right),\\ x^{*}=\arg\min_{x}\mathcal{L}\left(x,f_{\text{edge }}\left(x^{(k)}\right),z\right)\end{gathered}$ (2)
@@
According to [6], the loss function combines the $l_{2}$ distance between $f_{\text{edge }}\left(x^{(k)}\right)$ and $z$ with a TV regularization term for $x$, forming the basis of rMLE-based MIA. The method relies on gradient descent, where the loss values depend on output features, and gradients are computed using the target model’s parameters.
To achieve effective attacks, sufficient output feature information $\delta(z)$ and a well-chosen loss function are essential to ensure optimization space. Additionally, white-box knowledge of the edge model is required, as complex model mappings and limited feature information increase the risk of gradient descent converging to local minima [53, 11]. For instance, if two similar images produce nearly indistinguishable output features, reconstruction becomes challenging.
When dealing with complex edge model mappings or models protected by defense techniques, directly computing gradients in MLE-based MIA can significantly hinder optimization performance. To address this, He et al. [6, 16], building on the method from [10], developed a reverse mapping network for the edge model, which is categorized as Gen-based MIA. The key idea is to use a generator to approximate the reverse mapping of the edge model. Different methods adopt various strategies to train the generator. For instance, Yang et al. [10] trained a decoder using auxiliary task-related datasets, while Yin et al. [8] utilized intercepted feature data for training. The principle of this attack is summarized as follows.
$\begin{gathered}G=\arg\min_{G}\mathbb{E}_{x^{\prime}\sim p\left(x^{\prime}\right)}\mathcal{L}\left(G\left(f_{\text{edge }}\left(x^{\prime}\right)\right),x^{\prime}\right),\\ x^{*}=G(z)\end{gathered}$ (3)
@@
  1. 1.
    A well-matched training dataset aligned with the target task, ideally with a distribution similar to the target dataset [10].
  2. 2.
@@
The rationale is that if the information in $f_{\text{edge }}\left(x^{\prime}\right)$ is insufficient or cannot independently represent $x^{\prime}$’s features, the loss in (4) will struggle to converge as $G\left(f_{\text{edge }}\left(x^{\prime}\right)\right)$ remains ambiguous, limiting its training. Among them, the maximum separability is more decisive. For instance, even with random feature drops, if the remaining features distinguish inputs well, the attack can still reconstruct effectively [16]. Such frameworks resemble DAE [32], where the decoder can adapt to noise or feature drops.
@@ 4.2 Difficulty criterion for MIA in CI
This assumption holds primarily because the CI scenario involves deterministic inference, where model parameters are fixed and the feature extractor contains no stochastic components. Even if $x$ is perturbed by noise, the mapping $f_{\text{edge}}(x)$ still produces a unique and deterministic output $z$, ensuring that $H(z\mid x)=0$.
@@
These methods primarily compare $z$ with generated features, but a poorly trained or protected model can produce $z$ with high $\delta(z)$ yet low $H(z)$, indicating suboptimal separability. This weakens attack effectiveness despite the apparent information sufficiency. Thus, MLE-based MIA difficulty is more influenced by $I(x;z)$, $H(x|z)$, and $H(z)$.
To theoretically assess the difficulty of recovering the input $x$, we consider the reconstruction error probability $P_{e}$. In information theory, Fano’s inequality provides a fundamental lower bound on the reconstruction error in terms of conditional entropy. Modeling the MIA process as a Markov chain $x\to z\to x^{*}$, where $x^{*}$ is the reconstruction from $z$, Fano’s inequality [43, 42] impiles:
$H\left(P_{e}\right)+P_{e}\log|\mathbb{X}|\geq H(x\mid x^{*})\geq H(x\mid z)$ (5)
where $\mathbb{X}$ denotes the value space of $x$.
Since $|\mathbb{X}|\geq 2$ and $H(P_{e})\leq 1$, we can derive a simplified lower bound:
$P_{e}\geq\frac{H(x\mid z)-1}{\log|\mathbb{X}|}$ $=\frac{H(x)-I(x,z)-1}{\log|\mathbb{X}|}$ (6)
$=\frac{H(x)-H(z)-1}{\log|\mathbb{X}|}$
Here, $H(x)$ and $|\mathbb{X}|$ are considered intrinsic to the dataset and task, and typically remain fixed. Therefore, increasing the lower bound of $P_{e}$, and hence the difficulty of model inversion, is primarily influenced by $H(x\mid z)$, $H(z)$, and $I(x;z)$.
While Fano’s inequality traditionally assumes exact recovery of $x$, it remains applicable under relaxed semantic criteria, provided that the error probability $P_{e}$ is redefined accordingly. The inequality merely requires a binary distinction between success and failure, allowing $P_{e}$ to be interpreted in terms of semantic equivalence or perceptual similarity in the context of model inversion attacks.
Factors influencing Gen-based MIA. From Equation (4), Gen-based MIA requires sufficient input information $\delta(z)$ and high $H(z)$ to ensure maximum separability and enable loss convergence, allowing the generator to simulate reverse mapping effectively. A larger $\delta(z)$ not only facilitates generator training but also directly impacts the reconstruction of $x^{*}$. At the same time, the relationship in Inequality (7) also holds here [31]. Therefore, in addition to the factors influencing MLE-based MIA, Gen-based MIA is further effectively affected by $\delta(z)$.
@@
$D_{\text{mia}}\propto\left(\frac{k_{1}}{\{I(x;z);H(x\mid z);H(z)\}}+\frac{k_{2}}{\delta(z)}\right)$ (7)
Based on this criterion, existing defenses [21, 18, 34] using the information bottleneck theory can be understood as approximating a reduction in $I(x;z)$ through various algorithms to increase the difficulty of MIA. Perturbation-based defenses [16] aim to directly increase $H(x\mid z)$. The defensive effect of deeper neural networks [20, 15] can be explained by their ability to reduce $I(x;z)$ and $\delta(z)$. This is mainly because, without skip connections [44], mutual information progressively decreases during forward propagation in neural networks. In contrast, residual networks with skip connections preserve or even enhance $\delta(z)$ by adding layer outputs to inputs, but the decline in mutual information still poses challenges for MIA (see section 4.3 for experimental cases).
4.3 Demonstration and Guidance for Defense
To validate $D_{mia}$ and the theoretical analysis, we designed experiments using a CNN that progressively reduces the spatial size while increasing the channel size, based on the framework in [10, 38], and ResNet-18 with skip connections, quantifying each block. We trained MINE [24] on the same dataset used to train the full model $f$ to estimate the lower bound of $I(x,z)$. The core idea of MINE is to optimize a neural network estimator using the Donsker-Varadhan (DV) representation [45] as the loss function, providing an approximate estimate of the mutual information between random variables.
@@ 4.3 Demonstration and Guidance for Defense
$I(X;Z)=\mathbb{E}_{p(x,z)}\left[\log\frac{p(x,z)}{p(x)p(z)}\right]$ (8)
@@
$I(X;Z)\geq\sup_{T}\mathbb{E}_{p(x,z)}[T(x,z)]-\log\mathbb{E}_{p(x)p(z)}\left[e^{T(x,z)}\right]$ (9)
@@
$\hat{I}(X;Z)=\mathbb{E}_{p(x,z)}\left[T_{\theta}(x,z)\right]-\log\mathbb{E}_{p(x)p(z)}\left[e^{T_{\theta}(x,z)}\right]$ (10)
As shown in Figure 2, we trained the target models on the CIFAR-10 [46] and used Equation (11) to train MINE for the corresponding blocks. Figures 2(a, b) show that increasing network depth reduces both mutual information and effective information, leading to a decrease in attack performance. Figures 2(c, d) illustrate that while skip connections within two layers do not reduce $\delta(z)$, they slightly lower $I(x,z)$, moderately affecting attack performance. Previous research [44] identified vulnerabilities caused by skip connections in centralized MIA. Our study further highlights their impact on MIA effectiveness in CI.
Figure 2: The impact of neural network depth on $I(x,z)$, $\delta(z)$, as well as MLE & Gen-based MIA.
@@
Figure 3: Overview of SiftFunnel. Specifically, the design is divided into two aspects: the architecture and the loss function for the edge and cloud model.
Notably, while mutual information at ResNet’s fourth layer is higher than that at CNN’s second layer, MLE-based MIA does not perform better. This stems from the fact that these methods rely heavily on gradient descent, making them sensitive to suboptimal optimizers, hyper-parameter settings, and deeper network structures, which hinder gradient calculations for optimizing $x$. The lack of flexible and efficient methods for MLE-based MIA remains a gap in current research, often leading to their exclusion in defense evaluations—a topic warranting further exploration [20]. In our experiment and Section 6, we still evaluate MLE-based MIA against various defenses and provide evidence demonstrating the effectiveness of our method in defending against such attacks.
@@
5 SiftFunnel: Information Sifting Funnel
5.1 Overview of SiftFunnel
@@
Specifically, in the edge model design, the final layer adopts a funnel-shaped structure, reducing feature channels while preserving spatial dimensions. This reduces computation, lightens the edge device load, and limits redundant information. Attention mechanisms and multi-receptive field convolutions enhance task-relevant information, while a cloud-based upsampling module preserves accuracy. To reduce redundancy, $I(x,z)$ and increase the uncertainty of $z$ to $x$, we use Distance Correlation [47] for nonlinear correlations and the Pearson Coefficient [48] with $l_{1}$-norm for linear correlations. Additionally, Label Smoothing (LS) and Kullback-Leibler (KL) divergence [45] are used during training to improve model usability while reducing the separability of representations in $z$.
5.2 The Design of Shallow Edge Model
@@
Information Funnel. Deepening the neural network enhances feature extraction and reduces $I(x,z)$, as confirmed in Section 4. Previous work [49] shows through Fisher information [50] that deeper networks improve task relevance and resist MIA. Inspired by this, we redesign the final layer of the edge model with an Information Funnel module. This module reduces output channels to one-fourth in the first layer and further in the second layer, with channel selection based on the spatial size of feature maps. If the spatial size is reduced significantly, channels can exceed 10 to maintain accuracy. If unchanged, channels can be set to 2 to limit redundancy while preserving accuracy. Adjusting channels based on spatial size is necessary because spatial information loss is irreversible, whereas channel information loss can be compensated in the cloud model. This compensation increases $I(z,y)$ without reversing the reduction in $I(x,z)$, maintaining high attack difficulty as shown in Section 6.3. So, we add the channel compensation module in the cloud model to guarantee usability. Importantly, assumptions about extracting features from non-independent transmission modules are idealized, requiring adversaries to have full white-box knowledge of the cloud model as if they were the developer.
@@
The Information Funnel’s channel structure controls output but does not limit redundancy or enhance task-relevant information transmission, affecting usability. Inspired by Inception networks [51], we redesign it as a multi-receptive field convolutional framework with three low-channel layers. A $1*1$ kernel fuses channel information, followed by $3*3$ and $5*5$ kernels to expand the receptive field, extract richer task-relevant features, improve decision-making, and reduce mutual information with $x$, while lowering parameters of the edge model and controlling $\delta(z)$.
@@
Attention Block. To enhance the output of task-relevant information from the Information Funnel, attention modules are incorporated both before and after it. The first block combines an SE block [25] and Triplet Attention [26]. The SE block evaluates the contribution of information within each channel, suppressing channels with low task relevance. Triplet Attention reduces spatial information while maintaining smooth, continuous outputs, preserving fine-grained features for further processing. The second block integrates SE and CBAM modules [27]. Compared to Triplet Attention, CBAM is a lightweight module that sequentially calculates channel and spatial weights to enhance or suppress input features. Since CBAM’s channel and spatial attention are independently designed, certain weights are suppressed redundantly, resulting in sparser outputs. This sparsity is beneficial for output processing, effectively increasing $H(x\mid z)$ while reducing $H(z)$ and $\delta(z)$.
5.3 The Design of Loss Function